OpenStudy (anonymous):

Translate each phrase and select the one that is DIFFERENT from the others. the product of 2 and x twice x the value when x is doubled the square of x

OpenStudy (anonymous):

the product of 2 and x = 2*x =2x twice x = 2 times x = 2x the value when x is doubled = x+x = 2x the square of x =x*x = x^2 Now can you decide???

OpenStudy (anonymous):

C

OpenStudy (anonymous):

@Harkirat

OpenStudy (anonymous):

no it is the last one......it results in x^2 (i.e. square of x)